12 differently coloured beads are arranged around a necklace. How many different arrangements are possible?
Answer by ikleyn(52776)

.

This problem is about circular permutations.


The number of distinguishable circular permutations of "n" items is (n-1)!.


In this case the answer is (12-1)! = 11! = 11*10*9*8*7*6*5*4*3*2 = 39916800.


Solved.

-----------------

See the lesson
    - Persons sitting around a cicular table
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic  "Combinatorics: Combinations and permutations".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.